# Frontiers of Generative AI for Network Optimization: Theories, Limits, and Visions

### v1

@@ III-C GDMs for Other Network Tasks
While GDM has shown strong capabilities as a solution generator in combinatorial optimization and network optimization, its applications are not limited to these areas. GDM has also been extended to tackle black-box optimization (BBO) problems commonly found in AI4S scenarios [126]. Many BBO problems are leveraging GDM as a solution generator in materials science, molecular design [127, 128, 129], and topology optimization [130, 131, 127]. As shown in Fig. 5, BBO differs significantly from network optimization in both objectives and efficiency requirements. Specifically, the objective functions in BBO are unknown and cannot be expressed through explicit mathematical modeling, thus the only available feedback is the function value associated with a candidate solution, obtained through system evaluation. Moreover, BBO problems often have extremely large and complex solution spaces, which typically rely on large-scale random sampling of solution–value pairs to train models and use empirically low values to guide the generation process. For each distinct parameter configuration, a separate GDM-based generator must be constructed, which makes both training and inference inefficient, far from meeting the stringent efficiency demands of network optimization.
In addition, GDM continues to demonstrate innovation in conventional data synthesis tasks, particularly in areas such as channel estimation [132], channel denoising [133], and time-frequency sequence generation [134, 135].
@@ III-D Summary
As shown in Table V, both the solution generator and the DRL policy network of GDM in network optimization continue to evolve. However, the majority still focuses on GDM as a DRL policy. This trend primarily stems from the fact that many near-real-time resource allocation tasks in network optimization can be naturally formulated as MDP problems, making them well-suited for DRL-based approaches. Nevertheless, single-shot optimal optimization problems remain prevalent in the field, where GDM as a solution generator shows particular promise in enabling large-scale optimization and cross-problem generalization.
By comparing Table IV and Table V, an interesting contrast emerges regarding the diversity of neural backbones used in GDM-based solution generators and policies across domains. In the AI domain, GDM, as a solution generator, is mainly applied to classical combinatorial optimization problems, where graph convolutional networks (GCNs) dominate as the backbone architecture. In contrast, GDM as a policy network in this domain must address challenges such as multi-modal inputs and cross-task generalization, prompting exploration into a wider variety of neural architectures. In the networking domain, however, the situation is reversed. GDM-based solution generators are increasingly being designed for cross-problem generalization, leading to the adoption of new neural backbones. On the other hand, works that use GDM as a DRL policy typically focus more on novel scenario design and thus tend to adopt more established and stable network architectures for policy implementation.
Overall, GDM has contributed to a range of advances in network optimization, demonstrating effectiveness both as a solution generator for one-shot problems and as a policy network for MDP-based tasks. However, realizing its potential often requires meticulous problem-specific design. Despite the growing body of work, most existing research disregards the probabilistic nature of GDM, overlooking its inherent limitations when applied to tasks that demand precise mathematical reasoning and deterministic accuracy. Moreover, innovation in neural architectures and learning paradigms tailored for network optimization remains limited. To truly unlock GDM’s capabilities in this domain, future research should go beyond direct adaptation and explore architectures and training strategies that align more closely with the structural and precision demands of network optimization tasks.
Figure 5: Comparison between BBO and solution generator in the networking domain. Here, $\hat{y_{i}^{*}}$ denotes the hypothetical optimal solution for input $x_{i}$, while $\tilde{y_{i}^{*}}$ represents the corresponding generated solution.
@@ IV LPTMs in Network Optimization
TABLE V: Relevant works of GDM as a solution generator and DRL policy in network optimization.
{tblr}
colspec=X[0.14,l] X[0.22,c] X[0.15,c] X[0.63,c], cells = c, hlines, vlines, rowsep = 2pt, row1 = 0ex, cell21 = r=8, cell101 = r=14, Type & Reference Neural Backbone Contribution
@@
DiffSG [107] U-Net1D Demonstrated the broad applicability and cross-domain generalization potential of GDM.
@@
DRL policy GDMOPT [40] MLP Proposed a generalized implementation framework for GDM with DRL.
D2SAC [113] MLP Applied GDM with DRL to utility optimization in AIGC service systems.
Comparison of different GenAIs for DRL [114] MLP Demonstrated the superiority of GDM over other GenAI approaches as a DRL policy network in network optimization tasks.
SemSharing [115] MLP Applied GDM with DRL to incentive contract generation for information sharing in metaverse-VR environments.
G-DDPG [116] MLP Combined GDM with DDPG to optimize the denoising process in distributed diffusion model-based image generation.
DSAC-QMIX [117] MLP Applied GDM with DRL to optimize resource allocation and information freshness in tunnel construction wireless networks.
gMADRL-VCS [118] MLP Integrated GDM with multi-agent DRL to optimize post-disaster information collection in integrated space-air-ground networks.
GDM-based Multi-Dimensional Contract Design [119] MLP Formulated a contract theory and applied GDM with DRL to optimize incentive contracts for embodied vehicle intelligence in cooperative information sharing.
GDM-based Multi-Modal Offloading [120] Attention-based LSTM Applied GDM with DRL to parallel offloading and resource optimization in AIGC-metaverse-IoT systems.
Meta-DSAC [121] MLP Applied GDM with DRL to service provider selection optimization for AIGC-metaverse applications in UAV networks.
DLRS [122] MLP Used GDM with DRL to optimize cloud-edge collaborative deployment of LLMs.
GDM-driven Incentive Mechanism [123] MLP Applied GDM with DRL to resource allocation optimization for downstream tasks in semantic communication systems.
AC-GDM [124] MLP Applied GDM with DRL to IRS-assisted secure beamforming.
GAI-Asynchronous [125] MLP Applied GDM with DRL to batch-wise beamforming decision generation in low-altitude over-the-air edge inference systems.
@@
Specifically, TelecomGPT [141] addresses the lack of domain knowledge in general LLMs for telecom applications by building three datasets for pre-training, supervised fine-tuning, and style alignment. It targets tasks such as professional question-answering, document classification, code generation, code summarization, and mathematical modeling, achieving performance that surpasses general-purpose models, despite using models no larger than 8B parameters. Mobile LLaMA [142] focuses on network data analytics function and fine-tunes LLaMA 2 13B [70] on datasets covering IP routing analysis, packet inspection, and Python code diagnostics. M4 [143] proposes a firmware-like foundational model jointly managed by mobile operating systems and hardware, enabling support for dozens of multi-modal mobile AI tasks. M4 achieves accuracy comparable to baseline methods across most tasks while offering significantly reduced storage, memory, and operational complexity. WirelessGPT [144] targets channel-related tasks in wireless communications by introducing Traciverse, a large-scale wireless channel dataset for unsupervised pre-training. It encodes multi-dimensional channel data into unified representations while capturing inter-dimensional correlations. The model delivers strong performance in channel estimation, prediction, human activity recognition, and wireless signal reconstruction, with demonstrated scalability in model size. These foundational models developed for the networking domain provide valuable design insights and a rich set of benchmarks for applying LPTMs to network optimization tasks.
@@ IV-A LPTMs as Solution Generators
Although the natural language processing architecture of LLMs poses inherent limitations in directly solving network optimization problems (such as imprecision, high uncertainty, and significant computational overhead), several studies have nevertheless achieved meaningful results in direct solution generation by incorporating external feedback and prompt engineering. Specifically, [29] provides a forward-looking overview of generative optimization, suggesting that LLMs can offer hybrid improvements in three key areas: mitigating the computational cost of traditional optimization, enhancing initial solution quality through iterative exploration, and unifying inputs across different modalities. Furthermore, [83] proposes optimization by prompting (OPRO), which leverages an external objective evaluator to generate solution-score pairs. These historical evaluations are then used as prompts to iteratively guide the generation of new solutions. Similarly, [145] extends this approach to non-convex network optimization problems, demonstrating effective application beyond the simpler linear and low-dimensional cases considered in [83].
Overall, while LPTMs have demonstrated some effectiveness in directly generating solutions for network optimization through prompt engineering and iterative refinement, the computational overhead remains prohibitively high. Instead, their extensive pre-trained knowledge and strong cross-problem generalization capabilities hold greater promise as a key enabler for pushing the upper performance of GenAI-enabled network optimization.
@@ IV-B LPTMs as DRL Policy Networks
Overall, while LPTMs as DRL policy networks demonstrate impressive performance and often significantly outperform traditional architectures in terms of reward, they exhibit a poor performance-cost ratio, as the performance gains in network optimization do not justify the substantial training and computational overhead. Applying LPTMs to tasks such as intent decomposition and understanding, as well as high-level decision-making in time-insensitive scenarios, can be a more suitable direction.
@@ IV-C LPTMs for Other Network Tasks
At the current stage, LPTMs are still primarily used for understanding user intent [149, 116, 150, 151], designing and selecting customized network modules [152, 153], generating network algorithm code [154, 155, 156], and processing network data [157].
For user intent understanding, early work such as [149] proposes a shift from system-centric to user-centric paradigms. They build a pipeline of agents based on LLaMA-3B [70] to extract user intent and construct corresponding problem instances, which are then passed to external optimization solvers. [116] focuses on the distributed deployment of diffusion model-based image generation services. In this setting, LPTMs are used to construct cloned agents reflecting users’ personalized features, and the simulated feedback provided by these agents is leveraged to schedule the denoising process of the cloud-based diffusion models. LAMeTA [150] aims to enable intent-driven agentic networks. It distills a general-purpose LPTM into multiple edge-level LPTMs, each functioning as an agent within an agent graph. The core innovation lies in defining four theoretical metrics from the AI domain, including capability, information loss, latency, and error probability, to quantify service QoS requirements as 4-dimensional scalar vectors. These vectors are then used by an external DRL algorithm to plan the execution trajectory for each service request. T-ELLM [151] addresses challenges in federated learning caused by dynamically evolving network conditions. Based on natural language descriptions of service requirements, a fine-tuned LLM responds in text form with device selection results. External numerical solvers are then employed to carry out resource allocation.
TABLE VI: Relevant works of LPTMs in network optimization.
{tblr}
colspec=X[0.13,l] X[0.1,c] X[0.45,c] X[0.6,c], cells = c, hlines, vlines, rowsep = 2pt, row1 = 0ex, cell21 = r=3, cell51 = r=3, cell81 = r=4, cell121 = r=11, Type & Reference Characteristics Contribution
@@
[156] Network algorithm code generation Uses LLMs to generate hardware description language code on FPGA platforms.
@@
For customized network module design and selection, LLMind [152] proposes an intelligent framework to coordinate domain-specific AI modules. [153] introduces a framework that enhances interpretability in network synthesis by leveraging local sub-specifications corresponding to different network components. In particular, it incorporates constraint simplification techniques that, given configuration requirements and network topology, can generate local network configurations and assemble them into a complete setup.
For network algorithm code generation, [154] uses LPTMs to generate code for network management tasks and releases a benchmark and dataset named NeMoEval. NADA [155] performs block-wise code generation based on status prompts and neural architecture instructions. It then compiles and checks candidate code blocks, enabling flexible selection and assembly to meet evolving environmental demands. [156] leverages LLMs to generate hardware description language code for field programmable gate array (FPGA) platforms, using context learning and chain-of-thought prompting to achieve the desired functionality. The recent LLM-OptiRA [82] framework targets non-convex optimization in network scenarios by leveraging GPT-4 to transform non-convex problems into convex ones, subsequently generating executable code for mature numerical solvers to obtain precise solutions. LLM-OptiRA further incorporates code debugging and solution feasibility correction steps, achieving impressive code execution and solution success rates on the newly proposed WireOpt dataset.
For network data processing, Otter [157] presents an end-to-end solution compatible with various commercial embedded platforms and software. It processes data from embedded sensor terminals to edge devices, where LPTMs are used for prompting and key information filtering during preprocessing.
@@ IV-D Summary
As shown in Table VI, research on LPTMs in the network domain tends to focus on building foundation models for specific categories of tasks. Only subsequently is attention given to lower-level tasks related to network optimization, such as user intent understanding. Currently, the application of LPTMs in network optimization has evolved from initial efforts in multi-modal data processing and code generation to more advanced capabilities like user intent comprehension and complex task reasoning. This progression reflects a growing trend of adapting general-purpose LPTMs to a wide range of networking tasks, with flexible model sizes enabling practical deployment.
@@ V Theoretical Analysis of GenAI for Network Optimization
GenAI has demonstrated remarkable success in both one-shot solution generation and MDP-based decision-making for network optimization. However, there remains a lack of theoretical understanding regarding its generalization capabilities in optimization contexts. In this section, we take GDM as a representative example and apply the generalization capability analysis of distribution learning to the corresponding optimization problems. Specifically, we derive bounds on the gap between the expected objective value of generated solutions and the optimal value, as well as the gap between the expected returns of generated and optimal policies. These results provide a more rigorous foundation for understanding the generalization behavior of GenAI in both solution and decision generation, offering insights for practical applications.
@@ V-A Analysis
where $\hat{\boldsymbol{\theta}}_{n}$ represents the model parameters at training time $\tau$. Here, $m$ can be regarded as the model size, and $n$ is the number of training samples. $\bar{\mathcal{L}}$ and $\bar{\boldsymbol{\theta}}^{*}$ refer to the continuous version of the score function and its optimal parameterization, respectively, while $\mathcal{L}$ and $\boldsymbol{\theta}^{*}$ denote their generalized counterparts. $T$ is the total number of diffusion steps. $p_{T}$ and $\pi$ represent the noise distribution after $T$ steps of diffusion and the target noise distribution, respectively.
Both one-shot optimization and MDP decision-making in network optimization are modeled as conditional diffusion processes, which necessitates extending Eq. 1 to the conditional setting. Specifically, consider the conditional GDM here generation is conditioned on some variable $\mathbf{x}$ (input parameters), and let the joint distribution over solution and condition be denoted as $(\mathbf{y},\mathbf{x})$. We interpret the conditional model as learning the joint distribution $p(\mathbf{y},\mathbf{x})$, where the conditioning is implemented through distribution gradient guidance. Under this view, the true joint distribution can be formed as $p_{0}(\mathbf{y},\mathbf{x})$ and the model’s predicted joint distribution $p_{0|\hat{\boldsymbol{\theta}}_{n}(\tau)}(\mathbf{y},\mathbf{x})$. Then, the generalization bound in Eq. 1 naturally applies to the KL divergence $D_{\mathrm{KL}}(p_{0}(\mathbf{y},\mathbf{x})||p_{0|\hat{\boldsymbol{\theta}}_{n}(\tau)}(\mathbf{y},\mathbf{x}))$. Through the chain rule of KL divergence, we have
$D_{\mathrm{KL}}(p_{0}(\mathbf{y}$ $,\mathbf{x})||p_{0|\hat{\boldsymbol{\theta}}_{n}(\tau)}(\mathbf{y},\mathbf{x}))=D_{\mathrm{KL}}(p(\mathbf{x})||p_{\hat{\boldsymbol{\theta}}_{n}(\tau)}(\mathbf{x}))$ (2)
$+\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[D_{\mathrm{KL}}(p_{0}(\mathbf{y}|\mathbf{x})||p_{0|\hat{\boldsymbol{\theta}}_{n}(\tau)}(\mathbf{y}|\mathbf{x}))\right].$
In fact, the condition distribution $p(\mathbf{x})$ is preserved during training (e.g., drawn from the same distribution in both data and model), the first term vanishes or becomes negligible. In this case, the expected conditional KL divergence is bounded above by the joint divergence:
@@
This observation allows us to treat conditional diffusion models within the same generalization framework as unconditional ones, by simply extending the sample space to the joint variable $(\mathbf{y},\mathbf{x})$. Therefore, the theoretical result in Eq. 1 can be extended via Eq. 3 to analyze the generalization bound of conditional GDMs.
The following two subsections present how the KL-divergence-based generalization bound in Eq. 3 can be translated into generalization bounds for solution quality in optimization and expected return in decision-making, respectively.

### v2

@@ III-C GDMs for Other Network Tasks
While GDMs have shown strong capabilities as a solution generator in combinatorial optimization and network optimization, its applications are not limited to these areas. GDMs have also been extended to tackle black-box optimization (BBO) problems commonly found in AI4S scenarios [125]. Many BBO problems are leveraging GDMs as a solution generator in materials science, molecular design [126, 127, 128], and topology optimization [129, 130, 126]. As shown in Fig. 5, BBO differs significantly from network optimization in both objectives and efficiency requirements. Specifically, the objective functions in BBO are unknown and cannot be expressed through explicit mathematical modeling, thus the only available feedback is the function value associated with a candidate solution, obtained through system evaluation. Moreover, BBO problems often have extremely large and complex solution spaces, which typically rely on large-scale random sampling of solution–value pairs to train models and use empirically low values to guide the generation process. For each distinct parameter configuration, a separate GDM-based generator must be constructed, which makes both training and inference inefficient and falls short of meeting the stringent efficiency demands of network optimization.
In addition, GDMs continue to demonstrate innovation in conventional data synthesis tasks, particularly in areas such as channel estimation [131], channel denoising [132], and time-frequency sequence generation [133, 134].
@@ III-D Summary
As shown in Table VII, both the solution generator and the DRL policy network of GDMs in network optimization continue to evolve. However, the majority still focus on using GDMs as DRL policies. This trend primarily stems from the fact that many near-real-time resource allocation tasks in network optimization can be naturally formulated as MDP problems, making them well-suited for DRL-based approaches. Nevertheless, one-shot optimal optimization problems remain prevalent in the field, where using GDMs as solution generators shows particular promise in enabling large-scale optimization and cross-problem generalization.
By comparing Table VI and Table VII, an interesting contrast emerges regarding the diversity of neural backbones used in GDM-based solution generators and policies across domains. In the AI domain, GDMs, as solution generators, are mainly applied to classical combinatorial optimization problems, where graph convolutional networks (GCNs) dominate as the backbone architecture. In contrast, GDMs, as policy networks in this domain, must address challenges such as multi-modal inputs and cross-task generalization, prompting exploration of a wider variety of neural architectures. In the networking domain, however, the situation is reversed. GDM-based solution generators are increasingly being designed for cross-problem generalization, leading to the adoption of new neural backbones. On the other hand, works that use GDMs as DRL policies typically focus more on novel scenario design and thus tend to adopt more established and stable network architectures for policy implementation.
Overall, GDMs have contributed to a range of advances in network optimization, demonstrating effectiveness both as a solution generator for one-shot problems and as a policy network for MDP-based tasks. However, realizing its potential often requires meticulous problem-specific design. Despite the growing body of work, most existing research disregards the probabilistic nature of GDMs, overlooking its inherent limitations when applied to tasks that demand precise mathematical reasoning and deterministic accuracy. Moreover, innovation in neural architectures and learning paradigms tailored for network optimization remains limited. To truly unlock GDM’s capabilities in this domain, future research should go beyond direct adaptation and explore architectures and training strategies that better align with the structural and precision demands of network optimization tasks.
Figure 5: Comparison between BBO and solution generation in the networking domain. Here, $\mathbf{\hat{y_{i}^{*}}}$ denotes the hypothetical optimal solution for input $\mathbf{x}_{i}$, while $\mathbf{\tilde{y_{i}^{*}}}$ represents the corresponding generated solution.
@@ IV LPTMs in Network Optimization
TABLE VII: Relevant works of GDMs as solution generators and DRL policies in network optimization.
Type Reference Neural Backbone Contribution
@@
DiffSG [106] U-Net1D Demonstrated the broad applicability and cross-domain generalization potential of GDMs.
@@
DRL policy GDMOPT [40] MLP Proposed a generalized implementation framework for GDMs with DRL.
D2SAC [112] MLP Applied GDMs with DRL to utility optimization in AIGC service systems.
Comparison of different GenAIs for DRL [113] MLP Demonstrated the superiority of GDMs over other GenAI approaches as a DRL policy network in network optimization tasks.
SemSharing [114] MLP Applied GDMs with DRL to incentive contract generation for information sharing in metaverse-VR environments.
G-DDPG [115] MLP Combined GDMs with DDPG to optimize the denoising process in distributed diffusion model-based image generation.
DSAC-QMIX [116] MLP Applied GDMs with DRL to optimize resource allocation and information freshness in tunnel construction wireless networks.
gMADRL-VCS [117] MLP Integrated GDMs with multi-agent DRL to optimize post-disaster information collection in integrated space-air-ground networks.
GDM-based Multi-Dimensional Contract Design [118] MLP Formulated a contract theory and applied GDMs with DRL to optimize incentive contracts for embodied vehicle intelligence in cooperative information sharing.
GDM-based Multi-Modal Offloading [119] Attention-based LSTM Applied GDMs with DRL to parallel offloading and resource optimization in AIGC-metaverse-IoT systems.
Meta-DSAC [120] MLP Applied GDMs with DRL to service provider selection optimization for AIGC-metaverse applications in UAV networks.
DLRS [121] MLP Used GDMs with DRL to optimize cloud-edge collaborative deployment of LLMs.
GDM-driven Incentive Mechanism [122] MLP Applied GDMs with DRL to resource allocation optimization for downstream tasks in semantic communication systems.
AC-GDM [123] MLP Applied GDMs with DRL to IRS-assisted secure beamforming.
GAI-Asynchronous [124] MLP Applied GDMs with DRL to batch-wise beamforming decision generation in low-altitude over-the-air edge inference systems.
@@
Specifically, TelecomGPT [140] addresses the lack of domain knowledge in general LLMs for telecom applications by building three datasets for pre-training, supervised fine-tuning, and style alignment. It targets tasks such as professional question-answering, document classification, code generation, code summarization, and mathematical modeling, achieving performance that surpasses that of general-purpose models, despite using models with no more than 8B parameters. Mobile LLaMA [141] focuses on network data analytics function and fine-tunes LLaMA 2 13B [70] on datasets covering IP routing analysis, packet inspection, and Python code diagnostics. M4 [142] proposes a firmware-like foundational model jointly managed by mobile operating systems and hardware, enabling support for dozens of multi-modal mobile AI tasks. M4 achieves accuracy comparable to baseline methods across most tasks while offering significantly reduced storage, memory, and operational complexity. WirelessGPT [143] targets channel-related tasks in wireless communications by introducing Traciverse, a large-scale wireless channel dataset for unsupervised pre-training. It encodes multi-dimensional channel data into unified representations while capturing inter-dimensional correlations. The model delivers strong performance in channel estimation, prediction, human activity recognition, and wireless signal reconstruction, with demonstrated scalability in model size. These foundational models developed for the networking domain provide valuable design insights and a rich set of benchmarks for applying LPTMs to network optimization tasks.
@@ IV-A LPTMs as Solution Generators
Although the natural language processing architecture of LLMs poses inherent limitations for directly solving network optimization problems (such as imprecision, high uncertainty, and significant computational overhead), several studies have nevertheless achieved meaningful results by directly generating solutions through external feedback and prompt engineering. Specifically, [29] provides a forward-looking overview of generative optimization, suggesting that LLMs can offer hybrid improvements in three key areas: mitigating the computational cost of traditional optimization, enhancing initial solution quality through iterative exploration, and unifying inputs across different modalities. Furthermore, [82] proposes optimization by prompting (OPRO), which leverages an external objective evaluator to generate solution-score pairs. These historical evaluations are then used as prompts to iteratively guide the generation of new solutions. Similarly, [144] extends this approach to non-convex network optimization problems, demonstrating effective application beyond the simpler linear and low-dimensional cases considered in [82].
Overall, while LPTMs have demonstrated some effectiveness in directly generating solutions for network optimization through prompt engineering and iterative refinement, the computational overhead remains prohibitively high. Instead, their extensive pre-trained knowledge and strong cross-problem generalization capabilities hold greater promise as key enablers for improving the upper performance of GenAI-enabled network optimization.
@@ IV-B LPTMs as DRL Policy Networks
Overall, while LPTMs, as DRL policy networks, demonstrate impressive performance and often significantly outperform traditional architectures in terms of reward, they exhibit a poor performance-cost ratio, as the performance gains from network optimization do not justify the substantial training and computational overhead. Applying LPTMs to tasks such as intent decomposition and understanding, as well as high-level decision-making in time-insensitive scenarios, can be a more suitable direction.
@@ IV-C LPTMs for Other Network Tasks
At the current stage, LPTMs are still primarily used for understanding user intent [148, 115, 149, 150], designing and selecting customized network modules [151, 152], generating network algorithm code [54],[153, 154], and processing network data [155].
For user intent understanding, early work such as [148] proposes shifting from a system-centric to a user-centric paradigm. They build a pipeline of agents based on LLaMA-3B [70] to extract user intent and construct corresponding problem instances, which are then passed to external optimization solvers. [115] focuses on the distributed deployment of diffusion model-based image generation services. In this setting, LPTMs are used to construct cloned agents that reflect users’ personalized features, and the simulated feedback from these agents is leveraged to schedule the denoising process of cloud-based diffusion models. LAMeTA [149] aims to enable intent-driven agentic networks. It distills a general-purpose LPTM into multiple edge-level LPTMs, each functioning as an agent within an agent graph. The core innovation lies in defining four theoretical metrics from the AI domain, including capability, information loss, latency, and error probability, to quantify service QoS requirements as 4-dimensional scalar vectors. These vectors are then used by an external DRL algorithm to plan the execution trajectory for each service request. T-ELLM [150] addresses challenges in federated learning caused by dynamically evolving network conditions. Based on natural-language descriptions of service requirements, a fine-tuned LLM generates text responses with device selection results. External numerical solvers are then used to perform resource allocation.
TABLE VIII: Relevant works on LPTMs in network optimization.
Type Reference Characteristics Contribution
@@
For customized network module design and selection, LLMind [151] proposes an intelligent framework to coordinate domain-specific AI modules. [152] introduces a framework that enhances interpretability in network synthesis by leveraging local sub-specifications corresponding to different network components. In particular, it incorporates constraint-simplification techniques that, given configuration requirements and network topology, generate local network configurations and assemble them into a complete setup.
For network algorithm code generation, [153] uses LPTMs to generate code for network management tasks and releases a benchmark and dataset named NeMoEval. NADA [154] performs block-wise code generation based on status prompts and neural architecture instructions. It then compiles and checks candidate code blocks, enabling flexible selection and assembly to meet evolving environmental demands. The recent LLM-OptiRA [54] framework targets non-convex optimization in network scenarios by leveraging GPT-4 to transform non-convex problems into convex ones, subsequently generating executable code for mature numerical solvers to obtain precise solutions. LLM-OptiRA further incorporates code debugging and solution feasibility correction steps, achieving impressive code execution and solution success rates on the newly proposed WireOpt dataset.
For network data processing, Otter [155] presents an end-to-end solution compatible with various commercial embedded platforms and software. It processes data from embedded sensor terminals to edge devices, where LPTMs are used to prompt and filter key information during preprocessing.
@@ IV-D Summary
As shown in Table VIII, research on LPTMs in the network domain tends to focus on building foundation models for specific task categories. Only subsequently is attention given to lower-level tasks related to network optimization, such as user intent understanding. Currently, the application of LPTMs in network optimization has evolved from initial efforts in multimodal data processing and code generation to more advanced capabilities such as user intent comprehension and complex task reasoning. This progression reflects a growing trend of adapting general-purpose LPTMs to a wide range of networking tasks, with flexible model sizes enabling practical deployment.
@@ V Theoretical Analysis of GenAI for Network Optimization
GenAI has demonstrated remarkable success in both one-shot solution generation and MDP-based decision-making for network optimization. However, there remains a lack of theoretical understanding regarding its generalization capabilities in optimization contexts. In this section, we use GDMs as representative examples and apply the analysis of generalization capability in distribution learning to the corresponding optimization problems. Specifically, we derive bounds on the gap between the expected objective value of generated solutions and the optimal value, as well as the gap between the expected returns of generated and optimal policies. These results demonstrate the performance convergence of GDMs in both types of problem formulations and provide an intuitive illustration of the positive and negative factors influencing the convergence outcomes.
@@ V-A Analysis
where $\hat{\boldsymbol{\theta}}_{n}$ represents the model parameters at training time $\tau$. Here, $m$ denotes the model size, and $n$ denotes the number of training samples. $\bar{\mathcal{L}}$ and $\bar{\boldsymbol{\theta}}^{*}$ refer to the continuous version of the score function and its optimal parameterization, respectively, while $\mathcal{L}$ and $\boldsymbol{\theta}^{*}$ denote their generalized counterparts. $T$ is the total number of diffusion steps. $p_{T}$ and $\pi$ represent the noise distribution after $T$ steps of diffusion and the target noise distribution, respectively.
Both one-shot optimization and MDP decision-making in network optimization are modeled as conditional diffusion processes, which necessitates extending Eq. (1) to the conditional setting. Specifically, consider the conditional GDM, where generation is conditioned on some variable $\mathbf{x}$ (input parameters), and let the joint distribution over the solution and condition be denoted by $(\mathbf{y},\mathbf{x})$. Through the chain rule of KL divergence, the expected conditional KL divergence is bounded above by the joint divergence
@@
This observation allows us to treat conditional diffusion models within the same generalization framework as unconditional ones, by simply extending the sample space to the joint variable $(\mathbf{y},\mathbf{x})$.
Therefore, the convergence boundary of the GDM with respect to its original probabilistic distribution learning objective is established. Based on this boundary, the following two subsections derive performance bounds for the GDM in terms of one-shot optimization solution quality and MDP decision rewards, thereby characterizing the advantages and limitations of one-shot optimization and MDP decision-making within the probabilistic distribution-learning objective.